OpenStudy (anonymous):

(5x^(2/3))(4x^(1/4)) how do you multiply this

OpenStudy (anonymous):

20x^ (11/12)

jimthompson5910 (jim_thompson5910):

multiply the coefficients 5 and 4 to get 20, this is the final coefficient now add the exponents to get 2/3 + 1/4 = (2/3)*(4/4) + (1/4)*(3/3) = 8/12+3/12 =11/12 So the final exponent is 11/12 So the answer is 20x^(11/12)

OpenStudy (anonymous):

but what do you do with the exponenta

jimthompson5910 (jim_thompson5910):

you add them

OpenStudy (anonymous):

woahh i dont get what you did. do you need common denominator?

jimthompson5910 (jim_thompson5910):

sry, it's a bit jumbled 2/3 + 1/4 (2/3)*(4/4) + 1/4 ... Multiply top and bottom of the first fraction by 4 8/12 + 1/4 8/12+(1/4)(3/3) ... Multiply top and bottom of the second fraction by 3 8/12 + 3/12 (8+3)/12 11/12 So 2/3 + 1/4 = 11/12

jimthompson5910 (jim_thompson5910):

yeah, in order to add the fractions, you need to get all the denominators equal to the LCD (in this case 12)